OpenStudy (anonymous):

If an arrow is shot upwards on a certain planet with a velocity of 55m/s, its height in meters after t seconds is given by h =55t-0.04t^2. Find the average velocity of the arrow over the interval [1,1.04]. round to four decimials.

OpenStudy (anonymous):

I know tha tyou use the eqution (f(a-h) - f(x))/h

OpenStudy (anonymous):

PLEASE help

OpenStudy (dumbcow):

avg velocity = (h(1.04) -h(1))/(1.04-1) its just like finding slope

OpenStudy (anonymous):

that doesn't make sense how did you get that though

OpenStudy (dumbcow):

slope = (y2-y1)/(x2-x1) the calc definition yields same thing \[\frac{1}{b-a}\int\limits_{a}^{b}f'(x) dx = \frac{f(b)-f(a)}{b-a}\]